OpenStudy (anonymous):

Dirac delta function. Integral delta(t) from t1 to t2 for t1,t2<0

OpenStudy (anonymous):

I think it is 0 but am not sure

OpenStudy (anonymous):

Isn't it defined as the derivative of the Heaviside function?

OpenStudy (anonymous):

i think the dirac delta function is zero everywhere except at 0 so yes, the integral will be zero

OpenStudy (anonymous):

if i am not mistaken the integral \[\int _{-\infty}^{\infty}\delta=1\]

OpenStudy (anonymous):

some weird real tall function at \(x=0\) and \(0\) otherwise

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw, but i would not take my word for it maybe google

OpenStudy (anonymous):

I am sure now. I re read my notes and found it there too.

OpenStudy (anonymous):

whew

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."